OpenStudy (anonymous):

6t^2 - t - 1 / 3(t^2 + 1) = 2

OpenStudy (mathmate):

Cross multiply to get: 6t^2 - t - 1 = 2*3(t^2 + 1) Multiply out to get: t=-1-6=-7

OpenStudy (mathstudent55):